# Adapting Large Language Models for Character-based Augmentative and Alternative Communication

### v2

@@ 5.3 Sentence classification
Daily was the smallest with 46 K sentences. We used NLPAug Ma (2019) to add 232 K sentences, changing one or two words (selected at random) in the original sentences. Words were changed using a distilled version of RoBERTa Liu et al. (2019). This brought Daily to the same size (279 K sentences) as BOLT. We took the first 558 K sentences from Forum and News to create an equal amount of data for each of our three classes (in-domain written, in-domain spoken, and out-of-domain). During classification, we lowercased sentences and removed characters aside from A–Z, apostrophe, and space. We optimized hyperparameters as described in Section B.1. We used early stopping during training. Our final model had a classification accuracy of 85.7% on unseen sentences.
@@ 5.4 Domain adaptation procedure
We used the scored C4 and Subtitle sentences to adapt the LLM. We selected sentences with a written or spoken probability greater than or equal to a threshold. To efficiently determine which threshold to use, we trained 12-gram models from scratch across a broad range of thresholds and found we obtained the majority of perplexity gains with a threshold of 0.90 for C4 and 0.75 for Subtitle. Using these thresholds, our adaptation datasets consisted of 4.0 B words of C4 and 130 M words of Subtitle. With these thresholds as a starting point, we then trained opt-350m models with a narrower range of thresholds centered on 0.90 for C4 and 0.75 for Subtitle. The results, detailed in Section B.3, confirmed the thresholds we had found.
We first tested fine-tuning all the model’s parameters on a combined adaptation set consisting of C4 (threshold 0.90), Subtitle (threshold 0.75), BOLT, and Daily. As shown in Table 2, this reduced the perplexity on our dev sets from 2.30 to 2.19. Next, we explored a curriculum adaptation process consisting of a sequence of separate fine-tuning runs. Each run was on one of our five datasets: C4, Subtitle, Daily, BOLT, or Forum. Before each fine-tuning, we optimized hyperparameters using the model from the previous step.
Given the sizes of our five datasets, we opted to first always fine-tune on C4 followed by Subtitle. We tested all possible orders of further fine-tuning on Daily, BOLT, and Forum. Fine-tuning on Forum did not help in any order, and thus we excluded it from the curriculum. We found fine-tuning on C4, Subtitle, BOLT, and then Daily was the best order, improving perplexity to 2.10. Combining BOLT and Daily in a single final step did slightly better at 2.09. However, simply fine-tuning using the combined BOLT and Daily data did slightly better still at 2.07. While the classified C4 and Subtitle datasets did not prove useful for improving the subword LLM, they will be useful in improving other models (see Section 6).
@@
| C4+Sub+BOLT+Daily | full | 1 | 2.31 | 2.08 | 2.19 |
| C4SubBOLTDaily | full | 4 | 2.17 | 2.03 | 2.10 |
@@
| C4SubBOLT+Daily | LoRA | 3 | 2.18 | 2.06 | 2.12 |
@@
Additionally, we fine-tuned opt-350m using Low-Rank Adaptation (LoRA) Hu et al. (2021), which drastically reduces the number of trainable parameters for downstream tasks. As shown in Table 2 (bottom), LoRA yields slightly higher perplexities compared to full parameter fine-tuning. As such, we continued to use full fine-tuning for the remainder of our experiments.
For additional results and hyperparameter values, see Section B.2. We also explored other ways to try and improve adaptation. For details of things that did not work, see Section B.3.
@@ 6 Other prediction approaches
Our search algorithm (Section 3) takes on average 66 ms to predict the next character on a A100 GPU. While this may be fast enough for many purposes, interfaces such as Dasher Ward and MacKay (2002) or less capable devices may require more efficient prediction. We explored several other methods for obtaining faster predictions.
@@ 6.1 N-gram model
Inference with an n-gram model was very fast at 0.1 ms (Table 1). To serve as a baseline, and to provide an open training data n-gram model, we created a character 12-gram n-gram mixture model from C4 (threshold 0.90) and Subtitle (threshold 0.75) using SRILM Stolcke (2002). We created separate C4 and Subtitle language models using Witten-Bell smoothing and no count cutoffs. We built a mixture model using linear interpolation with weights optimized to minimize perplexity on dev sentences from Forum, BOLT, and Daily. The mixture weights were C4 0.84 and Subtitle 0.16.
The unpruned mixture model (denoted C4+Sub n-gram full) was large with 5 B parameters and a compressed size of 30 GB. We used entropy pruning Stolcke (1998) to to reduce the model to 71 M parameters and a compressed size of 488 MB (denoted C4+Sub n-gram). This was similar in size to the AAC n-gram we previously compared with (Table 1). On our dev sets, our full and pruned models had perplexities of 2.53 and 2.60 respectively. Our pruned model’s perplexity was somewhat higher than the 2.55 of the existing AAC n-gram model.
@@ 6.2 Byte model
The average perplexity on our dev sets of the base ByGPT5 model was 2.86. We compared our two best adaptation approaches found on opt-350m (Section 5.4). First, we used a three step curriculum approach, first adapting on C4, followed by Subtitle, and finally on BOLT combined with Daily. This yielded a model with a perplexity of 2.19.
Second, we fine-tuned in a single-step on BOLT combined with Daily. This results in a higher perplexity of 2.31. We suspect due to the multilingual nature of the base model, first fine-tuning on C4 and Subtitle helped move the model towards English. For results and hyperparameter values at each phase of the fine-tuning, see Section B.4.
| Language | Domain | Character | Written | Spoken | AAC user | ALS bank | COMM2 | Time |
| model | adaptation | prediction | (ppl) | (ppl) | (ppl) | (ppl) | (ppl) | (ms) |
| AAC n-gram | - | KenLM | 2.56 .03 | 2.51 .03 | 2.89 .01 | 2.33 .03 | 2.41 .02 | 0.1 |
| C4+Sub n-gram | - | KenLM | 2.62 .03 | 2.54 .02 | 2.86 .01 | 2.43 .03 | 2.45 .02 | 0.1 |
| C4+Sub n-gram full | - | KenLM | 2.57 .03 | 2.47 .02 | 2.82 .01 | 2.38 .03 | 2.40 .02 | 0.1 |
| opt-350m | - | search alg. | 2.40 .03 | 2.14 .02 | 2.24 .01 | 2.38 .03 | 2.20 .02 | 61.0 |
| opt-350m | C4SubBOLTDaily | search alg. | 2.15 .02 | 2.00 .02 | 2.29 .01 | 2.14 .03 | 2.01 .02 | 60.7 |
| opt-350m | C4SubBOLT+Daily | search alg. | 2.12 .02 | 2.00 .02 | 2.30 .01 | 2.14 .03 | 2.02 .02 | 60.6 |
| opt-350m | BOLT+Daily | search alg. | 2.10 .02 | 2.00 .02 | 2.29 .01 | 2.13 .03 | 2.01 .02 | 61.6 |
| ByGPT5 | C4SubBOLT+Daily | directly | 2.22 .03 | 2.11 .02 | 2.64 .01 | 2.22 .03 | 2.10 .03 | 11.9 |
| opt-350m | C4SubBOLT+Daily | classifier | 2.44 .03 | 2.48 .03 | 3.09 .02 | 2.41 .03 | 2.38 .02 | 15.3 |
@@ 6.3 Classification
We also explored making character predictions by adding a classification layer to the LLM. Instead of asking the model to generate the ensuing text given some context, we asked it to produce a distribution over classification labels – each label corresponded to a single character in our symbol set. We built the model on top of the best domain-adapted opt-350m model from Section 5.4. This allowed us to retain the model’s domain knowledge and focus additional training on the classification task.
@@
The classifier model did not perform as well as we had hoped, achieving a perplexity of 2.47. By comparison, training with BOLT and then Daily in separate steps yielded a slightly higher perplexity of 2.54. Training on just BOLT and Daily performed worse at 2.60. See Section B.5 for additional results and hyperparameter values.
@@ 7 Final evaluation
At this point, we have extensively used our private dev sets to guide our process. We conducted a final evaluation with our unseen private test sets and three additional test sets. Firstly, we created a test set based on the text of an AAC user. This set consisted of 3.6 K sentences (65 K words) written by a user of Dasher Ward and MacKay (2002). The text was from public and private presentations given by the user. We also added two public test sets: voice banked messages of people with ALS Costello (2014a), and the COMM2 set of responses to conversational situations Vertanen (2013).
@@
As shown in Table 3 (top), our C4+Sub n-gram model performed slightly worse with an average test set perplexity of 2.58 compared to the AAC n-gram model which averaged 2.54. Our unpruned full model did slightly better at 2.53. All n-gram models were outperformed by the LLMs, both with and without domain adaptation.
@@
As shown in Table 3 (middle), using the search algorithm with opt-350m adapted in a single step on the combination of BOLT and Daily performed the best on four of the five test sets. It tied with the three- and four-step curriculum method on the spoken test set (2.00) and with the four-step curriculum on the COMM2 set (2.01). The base opt-350m yielded the lowest perplexity on the test set from the AAC user (2.24), though the curriculum adapted opt-350m models were not far behind (2.29 – 2.30).
@@
Averaging the perplexity across all models, there was a noticeable increase in perplexity between our written (2.35) and spoken (2.25) test sets and the presentations given by the AAC user (2.60). The public test sets also produced lower average perplexities (2.28 for ALS bank and 2.22 for COMM2). While we hoped our data selection approach using diverse text from mobile forum messages would allow our adapted models to perform well on both simple everyday communications and more complex planned written communications, at least for this one user, this did not work. As a measure of text complexity, we calculated the number of characters per sentence. The AAC user’s sentences were on average 94.4 characters per sentence compared to the much shorter sentences in the written (31.4) and spoken (39.9) test sets.
@@
The byte LLM provided the fastest performance and was 5.1 times faster than our search algorithm and 1.3 times faster than using a classification layer.
@@ 8 Discussion
Most of our trained models had higher perplexities on the AAC test set than on any of the other test sets. The only exception was the base opt-350m model, which performed worst on the written test set. We believe this stems from the nature of the test sets. The written and spoken sets were created by crowdsourced workers writing about topics such as the weather or travel. As described in Section A.4, workers were instructed to invent questions or statements they might include in a conversation with another person. The ALS bank and COMM2 sets contain text of a similar nature. This aligns closely with the in-domain conversational datasets we used in our adaptation process.
@@
We obtained the lowest perplexities using a subword tokenized LLM and our search algorithm. While we investigated a multi-step curriculum to adapt the opt-350m model using sentences classified from large out-of-domain datasets, a simple fine-tuning using in-domain datasets worked better. The main drawback of the search algorithm is its inference time. If faster predictions are required for a particular interface or device, a byte or classification model may be preferred. However, currently subword LLMs are much more popular than byte LLMs. Further, the search approach can be used directly on an LLM without task specific training.
@@
It is also possible that the subword nature of the model prevented it from fully learning the character classification task. As discussed in Section 3, adding a new character can rearrange a word’s optimal tokenization. Because the model was originally trained to predict the next token in a sequence, it may not have considered all the different possible tokenizations as our search algorithm does.
@@ 9 Conclusions
We investigated how to leverage large transformer language models to aid non-speaking individuals using Augmentative and Alternative Communication (AAC) devices. In particular, we focused on how to generate the character predictions needed by interfaces where users write one letter-at-a-time. We presented an algorithm for obtaining character probabilities from LLMs using subword tokenization, and compared our algorithm to using a byte-level LLM or adding a classification layer on top of a subword LLM. We detail how we classified sentences in large corpora of web and movies subtitles text based on how similar sentences were to written and spoken style communications. We investigated a multi-step fine-tuning process using our scored sentences as well as our smaller in-domain corpora to adapt LLMs to our target domain.
We found that we were able to produce the lowest per-character perplexities using our search algorithm with the subword opt-350m model. We found that our domain adaptation curriculum was effective at improving the model performance for simple, conversational text, but did not improve performance for more complex prepared presentations from a single AAC user. Though it offered the best perplexity, our subword to character algorithm took over five times longer to produce character predictions than a native byte-level LLM.
To aide future research, we have made our unique datasets and models available.555https://osf.io/ajm7t/ This includes the model for classifying sentences as written or spoken, the scored C4 and Subtitle sentences, n-gram models trained on the classified sentences, fine-tuned opt-350m model, fine-tuned ByGPT5 model, and letter classification model.
@@
The ultimate goal of this work was to adapt large language models for use in AAC devices. With that goal in mind, it is a limitation that we did not have training data from actual AAC users. As we discussed in Section 4.1, such data is difficult to come by, so the datasets we created and used in this work aimed to approximate AAC-like text.
@@
The experiments we conducted as part of this research consisted solely of the offline evaluation of our models with fixed datasets. While the perplexity values we report are generally considered to be indicative of a model’s performance, they may not show the full extent of the impact on an actual user entering text. Text input is a complex task, especially when performed through an AAC device. Further research and user evaluation is required to fully understand the benefit or detriment of our tuned language models on the text entry performance of AAC users.
@@
Some recent work has explored freezing model layers that have relatively low training loss in order to focus fine-tuning on improving underperforming layers Yusufali et al. (2024). This was not a technique that we were able to explore in the scope of this work. However, we may have been able to improve the performance of the classification model by freezing the lower model layers that were already domain adapted and focusing the additional training on the classification layer specifically. We leave this experiment for future work.
This work focused on character predictions for AAC devices. However, even systems that operate on the character level may be able to benefit from word predictions and completions to further accelerate user input. Because models for word-level predictions have been explored previously, it was not something we studied in the scope of this work. We think it is likely that our domain adaptation process will benefit word predictions in addition to character predictions, but we did not conduct that evaluation.
@@
We only tested our search algorithm with a few model architectures. Of particular concern is our handling of token removal. Because the tokenizer used by the models in this work includes a space at the beginning of many of its tokens, our algorithm removes the end of the context up to and including the last space. Other models may have tokenizers that handle space differently. For example, the tokenizer for Mistral 7B v0.3 Jiang et al. (2023) ignores spaces between words. Our algorithm may need slight adjustments to work with such models.
During the development of our search algorithm, we conducted multiple rounds of performance optimization. We introduced various changes to the algorithm that, in the end, resulted in performance that was several orders of magnitude faster than our initial prototype. However, it is possible additional efforts could further improve performance. Further, currently our search algorithm is optimized for inference on a GPU and not a CPU. This could limit the AAC devices our algorithm can practically be deployed on.
@@
We primarily focused on fine-tuning approaches to domain adaptation and the task of classifying the next character. In-context learning (ICL) has emerged in recent years as an alternative for fine-tuning Dong et al. (2024). With in-context learning, a few examples of the task or domain are provided to the language model along with the query. The model takes the examples into account when producing its response, instead of requiring a separate training procedure. While ICL has shown promise, such emergent abilities may require large model sizes Wei et al. (2022). For privacy reasons, our focus here was on modest sized models that could be plausibly used and fine-tuned on an end user’s device. We were not able to explore the feasibility of ICL using modest sized models in the scope of this paper, and leave it for future work.
@@
Our fine-tuning process presented in this work relied heavily upon the Forum Vertanen and Kristensson (2021), BOLT Song et al. (2014), and Daily Li et al. (2017) datasets. While we did introduce additional larger datasets (C4 Raffel et al. (2020) and Subtitle Lison and Tiedemann (2016)), we filtered those sets to sentences that were similar to our in-domain sets. This achieved our intended result of improving the performance of our character prediction models. However, any biases present in Forum, BOLT, and Daily may also be reflected in the text selected from C4 and Subtitle. This could lead to predictions made by text entry interfaces relying on our models to skew towards those biases.
When developing AAC devices and interfaces that leverage language models to accelerate user input, it is crucial to support user autonomy. We developed our models to assist users in entering text based on the text that is most likely. However, there may be cases where users wish to enter text that is not likely, or that the model has not seen before. We encourage anyone that uses our models to consider this and include features in their interfaces and devices that allow the user to control when the model is applied.

### v3

@@ 5.3 Sentence classification
Daily was the smallest with 46 K sentences. We used NLPAug Ma (2019) to add 232 K sentences, changing one or two words (selected at random) in the original sentences. Words were changed using a distilled version of RoBERTa Liu et al. (2019). This brought Daily to the same size (279 K sentences) as BOLT. We took the first 558 K sentences from Forum and News to create an equal amount of data for each of our three classes (in-domain written, in-domain spoken, and out-of-domain). During classification, we lowercased sentences and removed characters aside from A–Z, apostrophe, and space. We optimized hyperparameters as described in Section˜B.1. We used early stopping during training. Our final model had a classification accuracy of 85.7% on unseen sentences.
@@ 5.4 Domain adaptation procedure
We scored each sentence in C4 and Subtitle using our classifier. To adapt the opt-350m LLM, we selected sentences with a written or spoken probability that met a threshold. To efficiently determine which threshold to use, we trained 12-gram models from scratch across a broad range of thresholds and found we obtained the majority of perplexity gains with a threshold of 0.90 for C4 and 0.75 for Subtitle. Using these thresholds, our adaptation datasets consisted of 4.0 B words of C4 and 130 M words of Subtitle. With these thresholds as a starting point, we then trained opt-350m models with a narrower range of thresholds centered on 0.90 for C4 and 0.75 for Subtitle. The results, detailed in Section˜B.3, confirmed the thresholds we chose.
We first tested fine-tuning all the model’s parameters on a combined adaptation set consisting of C4 (threshold 0.90), Subtitle (threshold 0.75), BOLT, and Daily. As shown in Table˜2, this reduced the perplexity on our dev sets from 2.30 to 2.19. Next, we explored a curriculum adaptation process consisting of a sequence of separate fine-tuning runs. Each run was on one of our five datasets: C4, Subtitle, Daily, BOLT, or Forum. Before each fine-tuning, we optimized hyperparameters using the model from the previous step.
Given the sizes of our five datasets, we opted to first always fine-tune on C4 followed by Subtitle. We tested all possible orders of further fine-tuning on Daily, BOLT, and Forum. Fine-tuning on Forum did not help in any order, and thus we excluded it from the curriculum. We found fine-tuning on C4, Subtitle, BOLT, and then Daily was the best order, improving perplexity to 2.10. Combining BOLT and Daily in a single final step did slightly better at 2.09. However, simply fine-tuning using the combined BOLT and Daily data did slightly better still at 2.07. While the classified C4 and Subtitle datasets did not prove useful for improving the subword LLM, they will be useful for creating or improving other models (see Section˜6).
@@
| C4+Sub+BOLT+Daily | full | 1 | 2.30 | 2.08 | 2.19 |
| C4SubBOLTDaily | full | 4 | 2.16 | 2.03 | 2.10 |
@@
| C4SubBOLT+Daily | LoRA | 3 | 2.17 | 2.06 | 2.11 |
@@
Additionally, we fine-tuned opt-350m using Low-Rank Adaptation (LoRA) Hu et al. (2021), which drastically reduces the number of trainable parameters for downstream tasks. As shown in Table˜2 (bottom), LoRA yields slightly higher perplexities compared to full parameter fine-tuning. As such, we continued to use full fine-tuning for the remainder of our experiments.
For additional results and hyperparameter values, see Section˜B.2. We also explored other ways to try and improve adaptation. For details of things that did not work, see Section˜B.3.
@@ 6 Other prediction approaches
With opt-350m, our search algorithm (Section˜3) takes on average 79 ms to predict the next character on an A100 GPU. While this may be fast enough for many purposes, interfaces such as Dasher Ward and MacKay (2002) or less capable devices may require more efficient prediction. We explored several other methods for obtaining faster predictions.
@@ 6.1 N-gram model
Inference with an n-gram model was very fast at 0.1 ms (Table˜1). To serve as a baseline, and to provide an open training data n-gram model, we created a character 12-gram n-gram mixture model from C4 (threshold 0.90) and Subtitle (threshold 0.75) using SRILM Stolcke (2002). We created separate C4 and Subtitle language models using Witten-Bell smoothing and no count cutoffs. We built a mixture model using linear interpolation with weights optimized to minimize perplexity on dev sentences from Forum, BOLT, and Daily. The mixture weights were 0.84 (C4) and 0.16 (Subtitle).
The unpruned mixture model (denoted C4+Sub n-gram full) was large with 5 B parameters and a compressed size of 30 GB. We used entropy pruning Stolcke (1998) to to reduce the model to 71 M parameters and a compressed size of 488 MB (denoted C4+Sub n-gram). This was similar in size to the AAC n-gram we previously compared with (Table˜1). On our dev sets, our full and pruned models had perplexities of 2.53 and 2.60 respectively. Our pruned model’s perplexity was somewhat higher than the 2.55 of the existing AAC n-gram model.
@@ 6.2 Byte model
The average perplexity on our dev sets of the base ByGPT5 model (289 M parameters) was 2.85. We compared our two best adaptation methods we previously found on opt-350m (Section˜5.4) using ByGPT5. First, we used a three step curriculum approach: adapting on C4, then on Subtitle, and finally on BOLT combined with Daily. This model had a perplexity of 2.14.
Second, we fine-tuned the base model in a single-step on BOLT combined with Daily. This results in a higher perplexity of 2.27. We suspect due to the multilingual nature of the base model, first fine-tuning on C4 and Subtitle helped move the model towards English. For results and hyperparameter values at each phase of the fine-tuning, see Section˜B.4.
| Language | Domain | Character | Written | Spoken | AAC user | ALS bank | COMM2 | Average | Time |
| model | adaptation | prediction | (ppl) | (ppl) | (ppl) | (ppl) | (ppl) | (ppl) | (ms) |
| AAC n-gram | - | KenLM | 2.56 | 2.51 | 2.89 | 2.33 | 2.41 | 2.54 | 0.1 |
| C4+Sub n-gram | - | KenLM | 2.62 | 2.54 | 2.86 | 2.43 | 2.45 | 2.58 | 0.1 |
| C4+Sub n-gram full | - | KenLM | 2.57 | 2.47 | 2.82 | 2.38 | 2.40 | 2.53 | 0.1 |
| opt-350m | - | search alg. | 2.40 | 2.15 | 2.24 | 2.39 | 2.20 | 2.28 | 79.1 |
| opt-350m | C4SubBOLTDaily | search alg. | 2.14 | 2.00 | 2.29 | 2.14 | 2.01 | 2.12 | 78.9 |
| opt-350m | C4SubBOLT+Daily | search alg. | 2.12 | 2.00 | 2.30 | 2.14 | 2.02 | 2.12 | 78.2 |
| opt-350m | BOLT+Daily | search alg. | 2.10 | 2.01 | 2.28 | 2.13 | 2.01 | 2.11 | 78.0 |
| ByGPT5 | C4SubBOLT+Daily | directly | 2.20 | 2.04 | 2.57 | 2.18 | 2.06 | 2.21 | 11.6 |
| opt-350m | C4SubBOLT+Daily | classifier | 2.42 | 2.42 | 3.03 | 2.38 | 2.36 | 2.52 | 14.6 |
| Language model | Domain adaptation | Written | Spoken | AAC user | ALS bank | COMM2 | Average |
| AAC n-gram | - | 59.7% | 60.0% | 55.3% | 62.2% | 61.0% | 59.7% |
| C4+Sub n-gram | - | 59.2% | 59.9% | 56.4% | 61.1% | 61.2% | 59.6% |
| C4+Sub n-gram full | - | 60.1% | 61.4% | 57.4% | 61.8% | 62.0% | 60.5% |
| opt-350m | - | 61.7% | 65.1% | 63.1% | 61.9% | 64.2% | 63.2% |
| opt-350m | C4SubBOLTDaily | 64.6% | 67.1% | 62.5% | 64.5% | 66.4% | 65.0% |
| opt-350m | C4SubBOLT+Daily | 64.9% | 67.0% | 62.3% | 64.4% | 66.3% | 65.0% |
| opt-350m | BOLT+Daily | 65.2% | 67.2% | 62.7% | 64.7% | 66.3% | 65.2% |
| ByGPT5 | C4SubBOLT+Daily | 64.3% | 66.8% | 61.6% | 63.7% | 65.9% | 64.5% |
@@ 6.3 Classification
We also explored making character predictions by adding a classification layer to the LLM. Instead of using the model to generate the ensuing text from some context, we asked it to produce a distribution over labels – each label represented a character. We built the model on top of the best domain-adapted opt-350m model from Section˜5.4. This retained the model’s domain knowledge and focused additional training on the classification task.
@@
The classifier model did not perform as well as we had hoped, achieving a perplexity of 2.47. By comparison, training with BOLT and then Daily in separate steps yielded a slightly higher perplexity of 2.54. Training on just BOLT and Daily performed worse at 2.60. See Section˜B.5 for additional results and hyperparameter values.
@@ 7 Final evaluation
At this point, we have extensively used our private dev sets to guide our process. We conducted a final evaluation with our unseen private test sets and three additional test sets. Firstly, we created a test set based on the text of an AAC user. This set consisted of 3.6 K sentences (65 K words) written by a user of Dasher Ward and MacKay (2002). The text was from public and private presentations given by the user. We also added two public test sets: voice banked messages of people with ALS Costello (2014), and the COMM2 set of responses to conversational situations Vertanen (2013).
@@
As shown in Table˜3 (top), our C4+Sub n-gram model performed slightly worse with an average test set perplexity of 2.58 compared to the AAC n-gram model which averaged 2.54. Our unpruned full model did slightly better at 2.53. All n-gram models were outperformed by the LLMs, both with and without domain adaptation.
@@
As shown in Table˜3 (middle), using the search algorithm with opt-350m adapted in a single step on the combination of BOLT and Daily performed the best on two of the five test sets and the overall average. It tied with the four-step curriculum on the COMM2 set (2.01). The base opt-350m yielded the lowest perplexity on the test set from the AAC user (2.24), though the curriculum adapted opt-350m models were not far behind (2.28–2.30).
@@
Averaging the perplexity across all models, there was a noticeable increase in perplexity between our written (2.35) and spoken (2.24) test sets and the presentations given by the AAC user (2.59). The public test sets also produced lower average perplexities (2.28 for ALS bank and 2.21 for COMM2). While we hoped our data selection approach using diverse text from mobile forum messages would allow our adapted models to perform well on both simple everyday communications and more complex planned written communications, at least for this one user, this did not work. As a measure of text complexity, we calculated the number of characters per sentence. The AAC user’s sentences were on average 94.4 characters per sentence compared to the much shorter sentences in the written (31.4) and spoken (39.9) test sets.
@@
Of the LLM approaches, the byte LLM provided the fastest predictions at around 12 ms. This was 6.8 times faster than our search algorithm and 1.3 times faster than using a classification layer.
To demonstrate the potential impact in a real-world AAC interface, we simulated using each language model in a onscreen keyboard with five word predictions. We simulated the keystrokes a hypothetical perfect user would save compared to typing every character. Keystroke savings is calculated as:
where is the number of keystrokes required without word predictions and is the number of keystrokes required with predictions. Higher keystroke savings is better.
The keyboard was allowed to make up to five word predictions. Predictions were made before any characters were typed for the next target word as well as during each subsequent prefix of the target word. The simulated user made no typing errors. We did not filter the offered word predictions based on whether the user did not choose a given prediction for a shorter prefix for the current target. If the target word appeared in the predictions, we assumed one key press selected the target word and added any following space.
As shown in Table˜4, keystroke savings improvements followed the perplexity gains (Table˜3). Compared to the AAC n-gram, across the test sets our best LLM provided a 5.5% absolute increase in keystroke savings. To contextualize this, imagine someone is using an eye-tracker and an onscreen keyboard with a dwell time of one second. Assuming they make no errors and incur no other overheads, writing the average length sentence in our spoken test set would take about 40 seconds without word predictions. Making optimal use of predictions provided by the AAC n-gram would take about 16.0 seconds. Using our best model (opt-350m adapted on BOLT+Daily), the user would take about 13.1 seconds, 18% faster than using the n-gram model for predictions.
@@ 8 Discussion
Most of our trained models had higher perplexities on the AAC test set than on any of the other test sets. The only exception was the base opt-350m model, which performed worst on the written test set. We believe this stems from the nature of the test sets. The written and spoken sets were created by crowdsourced workers writing about topics such as the weather or travel. As described in Section˜A.4, workers were instructed to invent questions or statements they might include in a conversation with another person. The ALS bank and COMM2 sets contain text of a similar nature. This aligns closely with the in-domain conversational datasets we used in our adaptation process.
@@
We obtained the lowest perplexities using a subword tokenized LLM and our search algorithm. While we investigated a multi-step curriculum to adapt the opt-350m model using sentences classified from large out-of-domain datasets, a simple fine-tuning using in-domain datasets worked better. The main drawback of the search algorithm is its inference time. If faster predictions are required for a particular interface or device, a byte or classification model may be preferred. However, currently subword LLMs are much more popular than byte LLMs. Further, the search approach can be used directly on an LLM without task-specific training.
@@
It is also possible that the subword nature of the model prevented it from fully learning the character classification task. As discussed in Section˜3, adding a new character can rearrange a word’s optimal tokenization. Because the model was originally trained to predict the next token in a sequence, it may not have considered all the different possible tokenizations as our search algorithm does.
@@ 9 Conclusions
We investigated how to leverage large transformer language models to aid non-speaking individuals using Augmentative and Alternative Communication (AAC) devices. In particular, we focused on how to generate the character predictions needed by interfaces where users write one letter-at-a-time. We presented an algorithm for obtaining character probabilities from LLMs using subword tokenization, and compared our algorithm to using a byte-level LLM or adding a classification layer on top of a subword LLM. We detail how we classified sentences in large corpora of web and movies subtitles text based on how similar sentences were to written and spoken communications. We investigated a multi-step fine-tuning process using our scored sentences as well as our smaller in-domain corpora to adapt LLMs to our target domain.
We found that we were able to produce the lowest per-character perplexities using our search algorithm with the subword opt-350m model. We found that our domain adaptation curriculum was effective at improving the model performance for simple, conversational text, but did not improve performance for more complex prepared presentations from a single AAC user. Though it offered the best perplexity, our subword to character algorithm took almost seven times longer to produce character predictions than a native byte-level LLM.
To aid future research, we have made our unique datasets and models available.555https://osf.io/ajm7t/ This includes the model for classifying sentences as written or spoken, scored C4 and Subtitle sentences, n-gram models trained on the classified sentences, fine-tuned opt-350m model, fine-tuned ByGPT5 model, and letter classification model.
@@
The ultimate goal of this work was to adapt large language models for use in AAC devices. With that goal in mind, it is a limitation that we did not have training data from actual AAC users. As we discussed in Section˜4.1, such data is difficult to come by, so the datasets we created and used in this work aimed to approximate AAC-like text. While we think our models can provide an improved initial experience for AAC text entry interfaces, we think much more substantial improvements are likely possible if models are adapted using the actual text of a given user. Doing this in a practical and privacy-preserving way will need further research.
@@
The experiments we conducted as part of this research consisted solely of the offline evaluation of our models with fixed datasets. While the perplexity values and keystroke savings we report are generally considered to be indicative of a model’s performance, they may not show the full extent of the impact on an actual user entering text. Text input is a complex task, especially when performed through an AAC device. Further research and user evaluation is required to fully understand the benefit or detriment of our tuned language models on the text entry performance of AAC users.
@@
Some recent work has explored freezing model layers that have relatively low training loss in order to focus fine-tuning on improving underperforming layers Yusufali et al. (2024). This was not a technique that we were able to explore in the scope of this work. However, we may have been able to improve the performance of the classification model by freezing the lower model layers that were already domain adapted and focusing the additional training on the classification layer specifically.
@@
We only tested our search algorithm with a few model architectures. Of particular concern is our handling of token removal. Because the tokenizer used by the models in this work includes a space at the beginning of many of its tokens, our algorithm removes the end of the context up to and including the last space. Other models may have tokenizers that handle space differently. Our algorithm may need adjustment to work with such models.
The main advantage of our search algorithm compared to the byte model is that it works out of the box with the most common LLMs which use subword tokenization. However, the main disadvantage is obtaining the distribution over the next character takes substantially longer. The search itself in many cases needs to make several inferences on the GPU so we can expect it to take several times as long as a byte LLM or classification head-based approach which can make the character prediction in a single inference. During the development of our search algorithm, we conducted multiple rounds of performance optimization. We introduced various changes to the algorithm that, in the end, resulted in performance that was several orders of magnitude faster than our initial prototype. However, it is possible additional efforts could further improve performance.
Currently, our search algorithm is optimized for inference on a GPU and not a CPU. This could limit the AAC devices our algorithm can practically be deployed on. However, for many AAC interfaces, such as those based on switch or brain input, waiting a tenth of a second for a prediction is likely acceptable and would not unduly impact user experience.
@@
We primarily focused on fine-tuning approaches to domain adaptation and the task of classifying the next character. In-context learning (ICL) has emerged in recent years as an alternative for fine-tuning Dong et al. (2024). With in-context learning, a few examples of the task or domain are provided to the language model along with the query. The model takes the examples into account when producing its response, instead of requiring a separate training procedure. While ICL has shown promise, such emergent abilities may require large model sizes Wei et al. (2022). For privacy reasons, our focus here was on modest sized models that could be plausibly used and fine-tuned on an end user’s device. We were not able to explore the feasibility of ICL using modest sized models in the scope of this paper, and leave this as future work.
@@
Our fine-tuning process presented in this work relied heavily upon the Forum Vertanen and Kristensson (2021), BOLT Song et al. (2014), and DailyDialog Li et al. (2017) datasets. While we did introduce the additional larger datasets C4 Raffel et al. (2020) and Subtitle Lison and Tiedemann (2016), we filtered those sets to sentences that were similar to our in-domain sets. This achieved our intended result of improving the performance of our character prediction models. However, any biases present in Forum, BOLT, and Daily may also be reflected in the text selected from C4 and Subtitle. This could lead to predictions made by text entry interfaces relying on our models to skew towards those biases.
When developing AAC interfaces that leverage language models to accelerate user input, it is crucial to support user autonomy, i.e. allowing users to express exactly what they want and not just what is probable under a language model. We developed our models to assist users in entering text based on the text that is most likely. However, there may be cases where users wish to enter text that is not likely, or that the model has not seen before. We encourage anyone who uses our models to design their interfaces such that users can control to what degree, if any, the language model contributes to the produced text.